OpenStudy (anonymous):

1/9-1/x over 1/81-1/x^2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

1 over 9 subtract 1 over x those two factors are then over 1 over 81 subtract 1 over x^2

OpenStudy (lgbasallote):

multiply everything by x^2 x^2/9 -x = x^2/81 -1 x^2/9 - x^2/81 - x = -1 729x^2/81 - x^2/81 - x = 81 728x^2/81 - x = 81 this is seriously complicated-___-

OpenStudy (anonymous):

i know right

OpenStudy (anonymous):

sorry

OpenStudy (lgbasallote):

sory that's 729x^2/81 -x = -1....not 81...maybe you'll have an idea

OpenStudy (anonymous):

answer choices are : x+9 over 9 or 9 over x+9 or x+9 over 9x or 9x over x+9

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

nevermind i got it

OpenStudy (anonymous):

1/9-1/x over 1/81-1/x^2 \[\huge\frac{\frac{1}{9}-\frac{1}{x}}{\frac{1}{81}-\frac{1}{x^2}}\]

OpenStudy (anonymous):

is this your problem?

OpenStudy (anonymous):

oh you got it

OpenStudy (anonymous):

yea thanks can u help on my next question

OpenStudy (anonymous):

i suppose, post it

OpenStudy (anonymous):

1/x-3+4/x/4/x-1/x-3

OpenStudy (anonymous):

you're going to have to use parentheses so i know what you are asking...

OpenStudy (anonymous):

1 over x-3 + 4 over x over 4/x - 1/x-3

OpenStudy (anonymous):

\[\huge\frac{\frac{1}{x-3}+\frac{4}{x}}{\frac{4}{x}-\frac{1}{x-3}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\huge\frac{\frac{x+4(x-3)}{x(x-3)}}{\frac{4(x-3)-x}{x(x-3)}}\]

OpenStudy (anonymous):

\[\huge{\frac{x+4(x-3)}{x(x-3)}}*\frac{x(x-3)}{4(x-3)-x}\]

OpenStudy (anonymous):

\[\frac{x+4(x-3)}{4(x-3)-x}\]

OpenStudy (anonymous):

\[\frac{x+4x-12}{4x-12-x}=\frac{5x-12}{3(x-4)}\]

OpenStudy (anonymous):

Do you understand how to do these?

OpenStudy (anonymous):

you must multiply each term in the numerator by the LCD of both terms, same thing with denominator

OpenStudy (anonymous):

i know how to setit up but i havent been able to find the answer

OpenStudy (anonymous):

The lcd of both the numerator and denominator are x(x-3), they will cancel because they divide by each other

OpenStudy (anonymous):

is the answer 5x^3/14

OpenStudy (anonymous):

no the answer is as i wrote it in my last post

OpenStudy (anonymous):

it can't be factored any further

OpenStudy (anonymous):

so the answrs 1 ?

OpenStudy (anonymous):

oo nvm

OpenStudy (anonymous):

thanks